Opened 8 years ago

Closed 7 years ago

Last modified 3 years ago

#368 closed flight processing (fixed)

EUFAR10/06 (ARMSRACE), flight day 231a/2010 Armsrace

Reported by: anch Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anch)

Data location: /users/rsg/arsf/workspace/EUFAR10_06-2010_231a_Armsrace

Data arrived from ARSF via SATA disk Q on 15/09/2010.

PI: Mate Stibranyi

Scientific objectives: N/A

Sensors:

  • Eagle (25/02/2011)
  • Hawk (25/02/2011)
  • Leica LIDAR (11/03/2011)
  • Digital Photography (11/03/2011)

Change History (28)

comment:1 Changed 8 years ago by anch

Please note that this project is stored in the archive-overflow directory until space becomes available.

comment:2 Changed 8 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:3 Changed 8 years ago by benj

  • Description modified (diff)

comment:4 Changed 8 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

Processed Applanix and IPAS

Starting Eagle and Camera processing

comment:5 Changed 8 years ago by emca

Also going to do the Hawk processing

comment:6 Changed 8 years ago by emca

Eagle and Hawk processing complete and ready for delivery check at
~airborne/workspace/EUFAR10_06-2010_231a_Armsrace/delivery/20110121

SCT values (and GPT) values for Eagle and Hawk lines:

Line no Eagle Hawk
1 -0.02 0.00
2 -0.06 (-0.09) -0.01
3 0.04 0.00
4 -0.13 (-0.14) 0.00
5 0.04 0.00
6 -0.03 -0.03
7 -0.02 0.00
8 0.03 0.00
9 -0.08 0.00
10 0.02 0.00

Notes:
Eagle/Hawk line 10 was processed using -l 400 8491 in azspec as it contained a bend.

Hawk raw files are numbered incorrectly (do not match the logsheet).
Raw files SWIR*10* correspond to line 1
Raw files SWIR*1* correspond to line 2
Raw files SWIR*2* correspond to line 3
.... etc

Lines were processed to be consistent with the logsheet and the Eagle line numbering.

comment:7 Changed 8 years ago by anch

Delivery checking hyperspectral

comment:8 Changed 8 years ago by anch

Delivery check complete.

Everything looks fine, although there is some cloud over the beginning and end of some lines, which could be mentioned in the data quality section. This is most obvious on lines 7, 8, and 9.

comment:9 Changed 8 years ago by anch

Starting LIDAR processing.

comment:10 Changed 8 years ago by emca

  • Description modified (diff)

Eagle and Hawk data sent to PI 25/02/2011

comment:11 Changed 8 years ago by anch

LiDAR submitted for delivery check

comment:12 Changed 8 years ago by emca

Camera ready for delivery check at
~airborne/workspace/EUFAR10_06-2010_231a_Armsrace/delivery/20110310/EUFAR10_06

comment:13 Changed 8 years ago by anch

Camera delivery checked.

All looks good an ready to deliver.

comment:14 Changed 8 years ago by anch

  • Description modified (diff)

LiDAR and Camera delivered (11/03/2011)

comment:15 Changed 8 years ago by emca

Delivery checked LiDAR delivery - no issues with data

comment:16 Changed 8 years ago by anch

  • Summary changed from EUFAR10/06, flight day 231a/2010 Armsrace to EUFAR10/06 (ARMSRACE), flight day 231a/2010 Armsrace

comment:17 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa

Project has been r-synced. Beginning archiving.

comment:18 Changed 8 years ago by knpa

Leica gps files are missing.

comment:19 Changed 8 years ago by knpa

tabi data present.

comment:20 Changed 8 years ago by knpa

There is some grimm data present.

comment:21 Changed 8 years ago by knpa

Comment above about ipas/gps is incorrect, it's the applanix/gps files which are missing.

comment:22 Changed 8 years ago by knpa

Projects has been tarballed and made available to NEODC (12/05/11)

comment:23 Changed 7 years ago by knpa

Backed-up by NEODC.

Moved to archive disk.

comment:24 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:25 Changed 7 years ago by emca

Reprocessed hyperspectral data with apl

New sct offsets for
Eagle 2: -0.03
Eagle 4: -0.07

comment:26 Changed 6 years ago by emca

Reprocessed the data for all bands and created delivery.

Ready for checking.

comment:27 Changed 6 years ago by anhi

Reprocessed hyperspectral delivery check - no issues.

Note: See TracTickets for help on using tickets.